OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 - 6x + 7 A. y = (x + 3)^2 + 2 B. y = (x - 3)^2 - 2 C. y = (x + 3)^2 - 2 D. y = (x - 3)^2 + 2

Parth (parthkohli):

Complete the square.

Parth (parthkohli):

\[\rm {\begin{aligned} y - 7&=x^ 2 - 6x \\ \rm y + 2&=x^2 -6x + 9 \\ \rm y + 2&=(x - 3)^2 \\ \rm y &=(x - 3)^2 - 2\end{aligned}}\]